OpenStudy (anonymous):

A 0.944 M solution of glucose,C6H12O6 , in water has a density of 1.0624g/ml at 20degree Celcius. What is the concentration of this solution in the following units? Mole fraction ? Mass Percent? Molality

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

mass of C6H12O6=72+12+96=180. M=moles/litre ,so moles of glucose per litre=0.944. mass per litre=0.944*180=169.92g total mass per litre = 1062.4g Hence mass of water=1062.4-169.92=894.28. moles of C6H12O6=0.944 mass of solvent in kg=0.89428 Molality is moles/mass of solvent = 0.944/0.894=1.055m

OpenStudy (anonymous):

Moles of water=894.28/18=49.68 Moles of glucose=0.944 Mole fraction=0.944/(0.944+49.68)=0.018. mass percent=169.92*100/(169.92+894.28)= 169.92*100/1062.4=0.15*100=15%.